### v1

# DESIGNER: Design-Logic-Guided Multidisciplinary Data Synthesis for LLM Reasoning

Abstract
Large language models (LLMs) have achieved remarkable success in many natural language tasks but still struggle with complex, multi-step reasoning, particularly across diverse disciplines. Existing reasoning datasets often either lack disciplinary breadth or the structural depth necessary to elicit robust reasoning behaviors. We propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages naturally available, extensive raw documents (book corpus and web corpus) to generate multidisciplinary challenging questions. A core innovation of our approach is the introduction of a Design Logic concept, which mimics the question-creation process of human educators. We use LLMs to reverse-engineer and abstract over 120,000 design logics from existing questions across various disciplines. By matching these design logics with disciplinary source materials, we are able to create reasoning questions that far surpass the difficulty and diversity of existing datasets. Based on this pipeline, we synthesized two large-scale reasoning datasets that span 75 disciplines: Design-Logic-Reasoning-Book (DLR-Book), containing 3.04 million challenging questions synthesized from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web), with 1.66 million challenging questions from the web corpus. Our data analysis demonstrates that the questions synthesized by our method exhibit substantially greater difficulty and diversity than those in the baseline datasets. We validate the effectiveness of these datasets by conducting SFT experiments on the Qwen3-8B-Base and Qwen3-4B-Base models. The results show that our dataset significantly outperforms existing multidisciplinary datasets of the same volume. Training with the full datasets further enables the models to surpass the multidisciplinary reasoning performance of the official Qwen3-8B and Qwen3-4B models. [^1] Project page: https://attention-is-all-i-need.github.io/Design-Logic-Reasoning
footnotetext: * First two authors contributed equally.    Corresponding Authors: Yongchi Zhao, Jiaheng Liu.
1 Introduction
Large language models (LLMs) have demonstrated exceptional capabilities in various natural reasoning tasks (Brown et al., 2020; Kaplan et al., 2020; Hoffmann et al., 2022; Chowdhery et al., 2023; OpenAI, 2023), such as mathematics and coding, especially when utilizing long chain-of-thought (CoT) techniques (Guo et al., 2025; Moshkov et al., 2025; Cai et al., 2025). However, their performance in a wide range of university-level discipline-specific reasoning questions still falls short of human experts. A major bottleneck is the lack of a large-scale, high-quality, and diverse multidisciplinary reasoning data. Existing datasets primarily focus on math and programming, training models on questions collected from math and coding competition platforms (Moshkov et al., 2025; Cai et al., 2025). This approach heavily relies on the massive open question resources in these specific domains, but many other academic disciplines lack such abundant resources, which limits the development of LLMs’ multidisciplinary reasoning capabilities.
Data synthesis is an effective way to address data scarcity by generating a large number of reasoning questions from existing LLMs (Wang et al., 2023; Xu et al., 2023; Jung et al., 2025; Havrilla et al., 2025). Current question synthesis methods can be broadly categorized into query-centric and doc-centric approaches. Query-centric methods, centered on an initial “seed” question pool, iteratively evolve existing questions. These methods generate more complex and diverse new questions by rewriting existing ones, adding constraints (e.g., Evol-Instruct), or incorporating a chain-of-thought (Wang et al., 2023; Xu et al., 2023; Yu et al., 2025). However, the breadth and depth of the generated questions are severely limited by the quality and coverage of the initial seed pool. They are also susceptible to the biases of the generative model itself, making it difficult to transcend domain limitations and create high-quality questions across a wide range of disciplines. In contrast, doc-centric methods start with a vast collection of unstructured (e.g., web pages, books) or structured (e.g., knowledge graphs) documents (Yue et al., 2024; Yuan et al., 2025; Huang et al., 2025). By directly extracting or reasoning out question-answer pairs from these documents, this approach ensures that the synthesized data is closely tied to specific domain knowledge and facts, theoretically allowing it to cover a broader range of disciplines. The main challenge for these methods is to effectively control the difficulty and diversity of the generated questions and prevent them from degrading into simple surface-level factual recall tasks.
To address these issues, as shown in Figure 1, we propose DESIGNER: a DESIGN-logic-guidEd Reasoning data synthesis pipeline that leverages readily available, large-scale, multidisciplinary raw documents (e.g., book corpus and web corpus) to synthesize a massive number of multidisciplinary challenging questions. The core innovation of our method lies in the introduction of a novel concept: “Design Logic”. We observed that when human education experts design challenging and insightful questions, they don’t merely state facts. Instead, they follow a structured design process, such as: identifying key knowledge points $\rightarrow$ constructing a scenario $\rightarrow$ designing a reasoning path $\rightarrow$ pre-setting distractors. We contend that this process embodies a “Design Logic”: a sequence of deliberate steps that transforms fundamental knowledge points into complex, context-rich questions that demand multi-stage reasoning. This design logic itself is a valuable and transferable form of “meta-knowledge”. Therefore, our pipeline is not limited to directly extracting questions from documents. Instead, we first use a powerful LLM to reverse-engineer and abstract tens of thousands of “meta design logics” from existing high-quality question banks across various disciplines. During the synthesis phase, we precisely match the most suitable design logic to a source document and instruct the LLM to strictly follow the intrinsic reasoning steps of that logic to construct a brand-new question.
Specifically, our approach follows a refined pipeline. First, we perform comprehensive processing on the massive book corpus and web corpus through multi-dimensional labeling and filtering (e.g., by discipline, readability, and knowledge value) to build a high-quality source material library that covers a wide range of disciplines. Next, from a private question bank containing hundreds of millions of questions, we use clustering and sampling to select a subset of questions that are both difficult and diverse. We then instruct a powerful LLM to summarize and abstract over 120,000 structured “design logics” from this subset. In the question synthesis phase, we pioneered a two-stage “retrieve-and-generate” matching mechanism. First, we use vector similarity to retrieve a rough, coarse-grained set of the most relevant candidate logics for each source document from our design logic library. Then, an LLM performs a fine-grained evaluation to select the optimal logic and strictly applies this design logic’s steps to the source document content, generating a new and challenging reasoning question.
The main contributions of this paper can be summarized as follows:
  • We propose a data engineering pipeline for synthesizing challenging questions from raw text corpora. Based on this pipeline, we constructed two large-scale reasoning datasets: Design-Logic-Reasoning-Book (DLR-Book), containing 3.04 million challenging questions synthesized from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web), with 1.66 million challenging questions from the web corpus. These datasets expand beyond common disciplines like math to cover 75 disciplines (including STEM, humanities, social sciences, arts, as well as applied and professional fields) to better train LLMs’ multidisciplinary reasoning capabilities.
  • We introduce a novel, scalable data synthesis paradigm based on “Design Logic”, which imitates the wisdom of human experts in question creation and significantly enhances the reasoning depth and diversity of doc-centric synthesized questions. Our data analysis demonstrates that the questions synthesized by our method exhibit substantially greater difficulty and diversity than those in the baseline datasets.
  • Our comprehensive comparative and ablation experiments on the Qwen3-8B-Base and Qwen3-4B-Base models (Yang et al., 2025a) show that training with our data not only significantly improves the model’s multidisciplinary reasoning capabilities but also outperforms existing multidisciplinary datasets with the same data volume. Training with our full dataset allows the resulting models to surpass the performance of the official Qwen3 models of the same size.
Figure 1: The Design-Logic-Guided Multidisciplinary Data Synthesis Pipeline.
2 Overview
We propose a data engineering pipeline for synthesizing challenging questions from raw text corpora (web corpus and book corpus). The complete pipeline is illustrated in Figure 1.
The first stage involves data curation, a meticulous process tailored to each of our three primary data sources and unified under a comprehensive 75-discipline taxonomy:
  • Web Corpus: We apply reasoning-oriented filtering and relabeling to a large web dataset, retaining only texts that are most useful for studying reasoning processes.
  • Book Corpus: This corpus undergoes preprocessing, discipline labeling, and a quality-prioritized sampling method to ensure the selection of high-readability and high-helpfulness text segments.
  • Question Bank: We annotate our question bank with multi-dimensional labels (discipline, difficulty, type) and use a clustering-based approach to select a diverse and high-quality subset.
The second stage focuses on question synthesis. We first extract underlying design logics from the sampled question subset. We then leverage a two-stage “retrieve-and-generate” mechanism to apply these reusable design logics to our curated text corpora. This process allows us to strictly follow the steps of each logic to construct millions of new reasoning questions, ensuring they are both challenging and diverse while mirroring the complexity of human-designed problems. Finally, we synthesize corresponding long CoT answers for each question, creating a valuable dataset for supervised fine-tuning.
Based on this pipeline, we constructed two large-scale reasoning datasets: Design-Logic-Reasoning-Book (DLR-Book), which contains 3.04 million challenging questions synthesized from the book corpus, and Design-Logic-Reasoning-Web (DLR-Web), with 1.66 million questions from the web corpus. These datasets expand beyond common disciplines like mathematics to cover 75 disciplines (including STEM and the humanities and social sciences), thereby better supporting the training of LLMs for multidisciplinary reasoning. Additionally, we have included concise reference answers in the datasets, generated from the text corpora concurrently with the questions, as well as long CoT responses generated by Qwen3-235B-A22B-Thinking-2507-FP8.
3 Data Curation
3.1 Data Collection
We curate three primary data sources and define a comprehensive discipline taxonomy:
  • Web Corpus: For the web corpus, we employ FineFineWeb [^2] https://huggingface.co/datasets/m-a-p/FineFineWeb, a filtered subset of the Common Crawl dataset.
  • Book Corpus: We utilize a proprietary library of books.
  • Question Bank: A proprietary repository of examination and practice items.
Discipline Taxonomy: We have established a classification system comprising 75 distinct disciplines, as shown in LABEL:tab:discipline_stats_en. This taxonomy provides comprehensive coverage across several major areas:
  • STEM: Science, Technology, Engineering, and Mathematics.
  • Humanities and Social Sciences: Including fields such as law, philosophy, and sociology.
  • Applied and Professional Fields: Encompassing domains like clinical medicine, education, and business administration.
  • Arts.
3.2 Data Processing and Filtering
3.2.1 Question Bank Processing
Multi-dimensional Labeling
Using Qwen3-30B-A3B (non-thinking), we annotate more than 150 million questions from the proprietary question bank with discipline, difficulty, and type labels. The specific prompts for these classifications are detailed in Figure 3, Figure 4, and Figure 5, respectively.
Clustering-Based Question Selection
To facilitate the extraction of design logic, we intend to filter a high-quality and diverse subset from the question bank. This subset is specifically curated to ensure a broad coverage of multiple disciplines and varying levels of difficulty. We compute embeddings with Qwen3-Embedding-4B and apply K-means clustering within each discipline (Ahmed et al., 2020). We then equally sample from each cluster with a difficulty ratio of Very Hard:Hard:Medium = 3:2:1, ensuring both diversity and sufficient difficulty. For each discipline, we determine the number of clusters via silhouette coefficient search and set per-discipline sample sizes according to the question bank’s discipline distribution. This process resulted in a curated set of 132,409 questions for design-logic extraction.
3.2.2 Book Corpus Processing
Text Preprocessing
We process each book at the chapter level. Chapters that exceed 5,000 words are further split into blocks up to 5,000 words. We then perform MinHash deduplication to obtain the final book segments.
Discipline Labeling and Quality Assessment
We assign discipline labels to book segments using a ModernBERT-large–based model that has been fine-tuned for this specific classification task (Warner et al., 2024). Subsequently, we introduce two key metrics for quality assessment: readability and helpfulness. Readability is predicted by a BERT-based classifier (Turc et al., 2019), which is designed to filter out incoherent or disorganized text. Helpfulness is scored by the fineweb-edu-classifier (Lozhkov et al., 2024), which provides a score on a scale of 0 to 5, allowing us to quantitatively assess the educational value of each document.
Quality-Prioritized Sampling
A total of over three million book segments were sampled from the corpus. For each discipline, we allocate quotas proportional to its frequency in both the book corpus and the question bank. To ensure the acquisition of high-quality data, we devised a quality-prioritized sampling policy. We first removed all samples for which the readability was classified as negative. Based on the helpfulness scores, the remaining candidate segments within each discipline are sorted in descending order, and samples are then selected from the top of this list until the predetermined quota is met. This process ensures that the sampled texts are of the highest quality available, with the vast majority of the final selected segments exhibiting a helpfulness score of $\geq 2$.
3.2.3 Web Corpus Processing
Reasoning-Oriented Filtering
Because not every FineFineWeb text contains sufficient reasoning content, we design a five-level scoring rubric (scores starting at 0; prompt in the Figure 6) to assess a text’s potential usefulness for studying reasoning processes. We utilized the Qwen3-30B-A3B (non-thinking) model to score 6.5 billion texts from the FineFineWeb dataset, retaining only those with a score of $\geq 3$.
Discipline Relabeling
Finally, we used the Qwen3-30B-A3B (non-thinking) model and the prompt specified in Figure 3 to relabel the retained texts, aligning the web corpus with our 75-discipline taxonomy.
4 Synthesizing Questions using Design Logic
4.1 Design Logic Extraction
Human educators typically construct exam questions by following a structured chain of design decisions that transforms simple knowledge concepts into complex and challenging questions rather than merely enumerating facts. A representative design process might involve: identifying the target objectives and knowledge points to be tested $\rightarrow$ constructing a contextual scenario $\rightarrow$ designing the logical reasoning process $\rightarrow$ formulating the correct answer $\rightarrow$ incorporating question traps and distracting information $\rightarrow$ validating the final question. Consequently, a solver must deconstruct the question through multi-step reasoning, which goes far beyond simple memorization (e.g., explaining a knowledge point).
Inspired by this practice, we propose a method for question synthesis based on “design logic”. Leveraging the prompt presented in Figure 7, we instruct an LLM (DeepSeek-R1-0528) to analyze the sampled authentic questions from our question bank. For each question, the LLM is tasked with three key steps: (i) deduce the question designer’s thought process, (ii) analyze how the question was constructed from the relevant knowledge points, and (iii) abstract and summarize the underlying design logic and principles. We specifically require the model to organize the resulting abstract design logic into the Mermaid format. This procedure yields a reusable pool of design logics, which the model can leverage these diverse logics to generate new questions based on provided source materials.
This approach eliminates the need for handcrafting a large set of disparate prompts, significantly increasing automation. It also substantially improves the difficulty and diversity of LLM-generated questions while aligning their structure more closely with questions authored by human teachers.
4.2 Design Logic Deduplication
To enhance the diversity of design principles provided to the LLMs, we further deduplicate the extracted design logic. This ensures a varied set of principles, which in turn facilitates the generation of diverse questions. Specifically, we use the Qwen3-Embedding-4B model to map each design logic into an embedding vector. For each discipline, we compute pairwise semantic similarities between all design logics to obtain a similarity matrix $S\in\mathbb{R}^{n\times n}$, where $S_{ij}$ denotes the semantic similarity between the $i$-th and $j$-th design logic.
Algorithm 1 Graph-based Deduplication via Centroid Selection
1:Input: A set of items $\mathcal{D}=\{d_{1},\dots,d_{n}\}$, a similarity matrix $S\in\mathbb{R}^{n\times n}$, a similarity threshold $\tau$.
2:Output: A deduplicated set of representative items $\mathcal{R}$.
3:
4:Initialize an undirected graph $G=(V,E)$ where $V=\{1,\dots,n\}$ and $E=\emptyset$.
5:Initialize the set of representatives $\mathcal{R}=\emptyset$.
6:
7:// Build a similarity graph where nodes are items and edges connect similar items.
8:for $i=1$ to $n$ do
9:  for $j=i+1$ to $n$ do
10:   if $S_{ij}>\tau$ then
11:     Add edge $(i,j)$ to $E$.
12:   end if
13:  end for
14:end for
15:
16:// Identify clusters of duplicates by finding connected components.
17:Let $\mathcal{C}\leftarrow\text{FindConnectedComponents}(G)$.
18:
19:// Select the most representative item (centroid) from each cluster.
20:for each connected component $C\in\mathcal{C}$ do
21:  Find centroid index $i^{*}=\underset{i\in C}{\arg\max}\sum_{j\in C,j\neq i}S_{ij}$.
22:  Add item $d_{i^{*}}$ to $\mathcal{R}$.
23:end for
24:
25:return $\mathcal{R}$.
We then apply graph-based connected-components deduplication. We model the $n$ design logics within a discipline as a graph $G=(V,E)$ with $|V|=n$. An undirected edge is added between nodes $i$ and $j$ if their semantic similarity $S_{ij}\geq\tau$. This construction forms connected components, where each component represents a group of mutually redundant items. From each connected component $C$, we select a unique representative — the item with the maximum sum of similarities to all other items in the component, and discard the rest. The complete procedure is detailed in Algorithm 1, with the similarity threshold set to $\tau=0.85$.
After deduplication, we retain a total of 125,328 unique design logics. The per-discipline counts are reported in Table LABEL:tab:discipline_stats_en.
4.3 Question Synthesis
To address the computational complexity of directly matching each design logic with every text segment, which would lead to a combinatorial explosion, we propose an approach similar to Retrieval-Augmented Generation (RAG). We compute the cosine similarity between the embeddings of each text segment and each design logic for each discipline-specific corpus (book or web) using the Qwen3-Embedding-4B model with a task-specific instruction (Figure 8). Specifically, for a text segment $t$ and a design logic $d$, we calculate $s(t,d)\,=\,\cos\!\bigl{(}\mathbf{e}(t),\,\mathbf{e}(d)\bigr{)}$. We then retain the top-5 design logics with the highest similarity scores as candidates for question synthesis for each text segment.
We then prompt an LLM (DeepSeek-R1-0528) to generate a challenging, graduate-level exam question (prompt detailed in Figure 9). Specifically, the LLM is instructed to (i) select the most suitable question-design logic from the top-5 candidates for the given source text, and (ii) strictly follow the corresponding logic and steps to construct a challenging question. This two-stage process acts as a coarse-to-fine ranking mechanism. The initial retrieval of the top-5 design logics using semantic similarity serves as a coarse-grained recall stage. The LLM’s subsequent selection of the most appropriate logic then functions as a fine-grained ranking stage. This strategy ensures a precise match between the design logic and the text, significantly enhancing the quality of the synthesized questions.
[…]
Question Deduplication and Decontamination
We implemented a two-stage filtering pipeline to manage redundancy and avoid evaluation leakage. First, we used MinHash-based detection to remove highly similar questions. Second, to address potential data contamination, we decontaminated our synthetic questions by comparing them against all evaluation benchmarks employed in this study. This was achieved through the standard 13-gram decontamination method, during which non-semantic punctuation was ignored. Any question flagged as similar is discarded.
Using the filtered book and web corpora from Section 3.2, we generate one corresponding reasoning question per text segment with DeepSeek-R1-0528. After deduplication and decontamination, our final reasoning dataset comprises 3,040,620 challenging questions synthesized from the book corpus and 1,658,541 challenging questions synthesized from the web corpus.
4.4 Response Synthesis
To demonstrate that our synthesized questions can effectively elicit and transfer the long CoT capabilities of a reasoning model and improve the performance of models trained on this data, we employ Qwen3-235B-A22B-Thinking-2507-FP8 to generate a corresponding long CoT response for each synthesized question. These question-response pairs are then used for supervised fine-tuning (SFT).
5 Data Analysis
To begin our quality assessment, we performed a quantitative analysis comparing the complexity and difficulty of our synthesized datasets (DLR-Book and DLR-Web) with the baseline datasets. This analysis focused on the distribution of question types, the lengths of questions and responses, the overall difficulty of the questions, and their diversity.
5.1 Question Type Analysis
As shown in Table 1, problem-solving questions form the majority of our synthesized datasets, constituting 64.92% of the book corpus and 63.91% of the web corpus. Multiple-choice questions are the next most common type, comprising 29.94% and 32.43%, respectively. This distribution analysis indicates that the composition of our synthesized questions is skewed towards reasoning-based types, as opposed to simple recall-oriented ones.
Table 1: Distribution of question types in our synthesized datasets.
Type DLR-Book DLR-Web
Problem-solving question 64.92% 63.91%
Multiple-choice question 29.94% 32.43%
Proof question 4.39% 2.67%
Other question types 0.75% 0.99%
5.2 Data Complexity
In Table 2, we present a comparison of the average lengths of questions and their corresponding responses for our two synthesized datasets against WebInstruct (Full) and NaturalReasoning. The responses were all generated by the Qwen3-235B-A22B-Thinking-2507-FP8 model. It is evident that the questions generated by our method, as well as the responses required to answer them, are significantly longer than those in WebInstruct (Full) (Yue et al., 2024) and NaturalReasoning (Yuan et al., 2025). This observation indirectly substantiates the high degree of difficulty and complexity inherent in our synthesized questions. This substantial increase in question length indicates that our synthesis method generates questions with more detailed context and more intricate premises, requiring a deeper understanding of the input. We likewise observe longer response lengths in our datasets, a direct consequence of the increased question complexity. This observation suggests that solving our synthesized questions requires a deeper chain of reasoning compared to the baselines.
Table 2: Average length (in characters) of questions and corresponding responses. The responses were all generated by the Qwen3-235B-A22B-Thinking-2507-FP8 model. The average lengths for WebInstruct (Full) and NaturalReasoning were calculated from a random sample of 304,181 instances.
Dataset Avg. Question Length Avg. Response Length
WebInstruct (Full) 180.43 11133.88
NaturalReasoning 332.08 17162.85
DLR-Book 1284.79 18090.06
DLR-Web 1205.96 17528.59
5.3 Difficulty Analysis
To provide a more direct measure of difficulty, we used the Qwen3-30B-A3B-Instruct-2507 model to assign difficulty labels to questions from our book and web corpora, as well as from WebInstruct (Full) and NaturalReasoning. Based on the complexity and length of the required reasoning, we categorized questions into four difficulty levels: Easy, Medium, Hard, and Very Hard.
As shown in Table 3, our datasets are significantly more difficult. The combined percentage of “Hard” and “Very Hard” questions is notably high: 89.84% for DLR-Book and 83.15% for DLR-Web. This contrasts sharply with the baseline datasets. DLR-Book contains 54.66% “Very Hard” questions, substantially more than the 31.11% in NaturalReasoning and 3.59% in WebInstruct (Full). Conversely, the proportion of “Easy” questions in our datasets is negligible (0.27% and 0.72%), especially when compared to WebInstruct (Full)’s 39.02%.
Collectively, these quantitative analyses of question length and difficulty distribution consistently demonstrate that our data synthesis method produces questions of significantly greater complexity and difficulty than the baseline datasets.
Table 3: Difficulty distribution of questions across datasets.
Dataset Easy Medium Hard Very Hard
WebInstruct (Full) 39.02% 39.58% 17.84% 3.59%
NaturalReasoning 2.10% 26.25% 40.54% 31.11%
DLR-Book 0.27% 9.88% 35.18% 54.66%
DLR-Web 0.72% 16.13% 36.24% 46.91%
5.4 Diversity Analysis
To evaluate the semantic diversity of the synthesized questions, we conducted a quantitative analysis using distance-based metrics. This approach assesses diversity by examining the distribution of question embeddings in a high-dimensional semantic space.
Diversity Metrics and Experimental Setup
We first generated high-dimensional vector representations for each question using the Qwen3-Embedding-4B model. For a given question set $X=\{x_{1},x_{2},\dots,x_{N}\}$, this process yields a corresponding set of embedding vectors $E=\{e_{1},e_{2},\dots,e_{N}\}$, where $e_{i}\in\mathbb{R}^{d}$ is the $d$-dimensional embedding for question $x_{i}$. We uniformly sample N = 300,000 questions from each dataset and compute the following five distance-based metrics in the embedding space.
  1. 1.
    Mean Cosine Distance (DistMean Cosine) (Yang et al., 2025b): The average cosine distance between all unique pairs of embeddings, calculated as $M_{\text{cosine}}=\frac{2}{N(N-1)}\sum_{i<j}(1-\cos(e_{i},e_{j}))$. A higher value indicates greater semantic dissimilarity.
  2. 2.
    Mean L2 Distance (DistMean L2) (Yang et al., 2025b): The average Euclidean distance between all unique pairs of embeddings, calculated as $M_{\text{L2}}=\frac{2}{N(N-1)}\sum_{i<j}\|e_{i}-e_{j}\|_{2}$. This measures the average separation of questions in the embedding space.
  3. 3.
    1-Nearest Neighbor Distance (1-NN Distance) (Stasaski & Hearst, 2022): The average cosine distance from each embedding to its single nearest neighbor, given by $M_{\text{1-NN}}=\frac{1}{N}\sum_{i}d_{1}(e_{i})$, where $d_{1}(e_{i})$ is the cosine distance from $e_{i}$ to its closest neighbor. This metric highlights the presence of tightly clustered, near-identical questions.
  4. 4.
    Cluster Inertia (Du & Black, 2019): The total squared distance of samples to their closest cluster center after applying the K-means clustering algorithm. It is calculated as $M_{\text{inertia}}=\sum_{i=1}^{N}\min_{j}\|e_{i}-c_{j}\|_{2}^{2}$, where $c_{j}$ are the cluster centroids. This measures the overall spread and density of the data clusters.
  5. 5.
    Radius (Lai et al., 2020): The geometric mean of the standard deviations of the embedding dimensions, modeling the data as a multi-dimensional Gaussian distribution: $M_{\text{Radius}}=(\prod_{j=1}^{d}\sigma_{j})^{1/d}$, where $\sigma_{j}$ is the standard deviation along the $j$-th dimension. It directly quantifies the spread of the data in the semantic space.
As detailed in Table 4, our datasets, DLR-Book and DLR-Web, consistently demonstrate a greater diversity of questions compared to the baseline datasets across all five semantic diversity metrics. The higher Mean Cosine Distance and Mean L2 Distance values confirm that our synthesized questions are, on average, more semantically distinct than those in WebInstruct (Full) and NaturalReasoning, indicating a broader conceptual scope. The most notable difference is observed in the 1-NN Distance, where our datasets score approximately twice as high as the baselines. This suggests our method generates far fewer semantically redundant questions. Furthermore, the Cluster Inertia and Radius scores for our datasets indicate that the generated questions occupy a larger and more varied volume within the semantic embedding space. This quantitative evidence confirms that our synthesis pipeline produces not only more complex and difficult questions but also a significantly more diverse set of questions.
Table 4: Semantic diversity metrics for our synthesized datasets and the baseline datasets, calculated on a uniform sample of 300,000 questions from each. Higher is better for all metrics shown.
Dataset DistMean Cosine DistMean L2 1-NN Distance Cluster Inertia Radius
WebInstruct (Full) 0.7762 1.2436 0.1830 205,590.02 0.0169
NaturalReasoning 0.8233 1.2818 0.1915 226,288.91 0.0173
DLR-Book 0.8471 1.3008 0.3726 238,100.2656 0.0176
DLR-Web 0.8494 1.3026 0.3897 238,039.5000 0.0177
6 Experiments
In this study, we evaluate the effectiveness of our synthesized data by conducting supervised fine-tuning (SFT) on base models. The resulting models are then assessed on a diverse range of widely-used benchmarks. Our extensive experiments demonstrate that the data synthesized by our method can effectively elicit the long CoT reasoning capabilities of LLMs for complex, multidisciplinary questions. Furthermore, with the same amount of training data, our method yields superior model performance compared to existing datasets like NaturalReasoning and WebInstruct (Full).
6.1 Experimental Setup
SFT Settings
In the majority of our experiments, we perform SFT on the Qwen3-8B-Base model. Subsequently, we also validate the effectiveness of our synthesized data across other model scales. For all SFT experiments, we adhere to the hyperparameters detailed in Table 5.
Table 5: Hyperparameters for Supervised Fine-Tuning (SFT).
Parameter Value
Epoch 6
Batch Size 64
Learning Rate 1e-5
Learning Rate Schedule Cosine decay to 0
Evaluation Settings
To ensure a fair comparison, we employ a zero-shot evaluation setting for all trained models, using the consistent generation configuration specified in Table 6. For each benchmark, we conduct $N$ independent sampling rollouts for each test instance. Our evaluation involves the following three metrics:
  • Pass@1: The mean accuracy (%) across the $N$ rollouts. We report both the Pass@1 score and its standard deviation to show the performance and stability of our method.
  • CoT-SC: The accuracy (%) determined by a majority vote over the $N$ generated samples. This metric is equivalent to Self-Consistency with Chain-of-Thought (CoT-SC) method and is reported for all cases where $N>1$.
  • Pass@N: The proportion (%) of questions for which at least one correct solution is found among $N$ rollouts.
Table 6: Generation configuration for evaluation.
Parameter Value
Temperature 0.6
Top-K 20
Top-P 0.95
Max Context Length 32768
Benchmarks. The benchmarks used in our evaluation, along with their respective disciplines and the number of rollouts ($N$), are detailed in Table 7.
Table 7: Evaluation benchmarks, their disciplines, and the number of rollouts (NN).
Benchmark Disciplines Rollout ($N$)
AIME 2024 Mathematics 64
AIME 2025 Mathematics 64
MATH-500 (Lightman et al., 2024) Mathematics 4
GPQA-Diamond (Rein et al., 2024) Physics, Chemistry, Biology 10
GPQA-Main (Rein et al., 2024) Physics, Chemistry, Biology 10
SuperGPQA (Du et al., 2025) 285 graduate-level disciplines across 13 fields (e.g., Engineering, Management, Economics, Education, History, Science, Medicine, Law, Sociology, Philosophy, Agriculture, Literature) 1
MMLU (Hendrycks et al., 2020) 57 disciplines including STEM, humanities, social sciences, and professional fields 1
MMLU-Pro (Wang et al., 2024) Math, Physics, Chemistry, Law, Engineering, Economics, Health, Psychology, Business, Biology, Philosophy, Computer Science, History, Other 1
6.2 Supervised Fine-Tuning Experiments
We performed supervised fine-tuning (SFT) on the Qwen3-4B-Base [^3] https://huggingface.co/Qwen/Qwen3-4B-Base and Qwen3-8B-Base [^4] https://huggingface.co/Qwen/Qwen3-8B-Base models using our synthetic datasets, DLR-Book and DLR-Web. The resulting SFT models were then compared against the official Qwen3-4B [^5] https://huggingface.co/Qwen/Qwen3-4B and Qwen3-8B [^6] https://huggingface.co/Qwen/Qwen3-8B models (thinking mode) under our evaluation settings.
Table 8: Performance of models fine-tuned on our datasets. The models are trained on DLR-Web, DLR-Book, or the combined DLR-Web+Book dataset. Baselines include the official Qwen3-4B and Qwen3-8B models (thinking mode), evaluated under our settings. The best and second-best scores in each column are shown in bold and underlined, respectively.
Model MMLU MMLU-Pro GPQA-Diamond GPQA-Main SuperGPQA
Pass@1 Pass@1 Pass@1 CoT-SC Pass@1 CoT-SC Pass@1
[…]
As shown in Table 8, supervised fine-tuning with our synthetic datasets significantly improves model performance. For both the 4B and 8B model scales, our fine-tuned models consistently outperform the official Qwen3-4B and Qwen3-8B (thinking mode) baselines across all reasoning benchmarks. Notably, models trained on the combined “DLR-Web+Book” dataset achieve the best results across all benchmarks, while those trained on the “DLR-Book” dataset achieve the second-best. This improvement is particularly pronounced on highly complex reasoning tasks like GPQA-Diamond. These results affirm the efficacy of our data synthesis strategy for enhancing the reasoning capabilities of language models.
6.3 Performance on Advanced Mathematical Reasoning
For challenging mathematical competition questions, such as those from AIME, we additionally report Pass@N results. Prior work has shown that methods like Reinforcement Learning with Verifiable Rewards (RLVR) primarily enhance Pass@1 accuracy by better selecting among pre-existing reasoning paths, without significantly improving Pass@N rates (Yue et al., 2025). While our models’ comparatively lower Pass@1 scores are expected due to our exclusive use of SFT without a subsequent RL phase, the significant improvement in Pass@N for our fine-tuned models is noteworthy. This result suggests that our data successfully introduces novel reasoning patterns, thereby expanding the model’s solution space and increasing its problem-solving potential. This expanded capability provides a strong foundation for future work to apply RL and convert these reasoning skills into higher Pass@1 performance.
Table 9: Performance on mathematical reasoning benchmarks, including Pass@N results. The best results in each column are in bold, and the second-best are underlined. The models trained on DLR-Web+Book show marked improvements in Pass@N, indicating an expanded reasoning capability.
Model MATH-500 AIME 2024 AIME 2025
Pass@1 CoT-SC Pass@N Pass@1 CoT-SC Pass@N Pass@1 CoT-SC Pass@N
[…]
6.4 Data Scaling Effects
To validate the scalability of our data synthesis methodology, we conducted a series of experiments to evaluate the reasoning capabilities of models trained on progressively larger datasets synthesized by our approach. The results, illustrated in Figure 2, demonstrate a clear and positive correlation between the volume of synthetic data and model performance across multiple challenging reasoning benchmarks.
The consistent improvement across this diverse set of benchmarks validates that our synthesis process is not overfit to a specific domain or benchmark but rather enhances a general, robust reasoning capability. These strong scaling laws confirm that our method provides a reliable pathway to achieving superior model performance through data augmentation, and future work can leverage our proposed design logics to synthesize even larger datasets for continued improvement.
Figure 2: The model’s performance on four key benchmarks (AIME 2025, MMLU-Pro, GPQA-Diamond, and SuperGPQA) as a function of the training data quantity. The performance metric is Pass@N for AIME 2025, while all others are Pass@1.
6.5 Comparison with Baseline Datasets
We conducted a comparative analysis between our synthetically generated data and other prominent open-source synthetic datasets. To ensure a fair comparison, given the varying sizes of the original datasets, we randomly sampled an equal number of 304,181 instances from each dataset for the fine-tuning experiments. For the purpose of equitable comparison, we regenerated long CoT responses for all baseline methods (including WebInstruct (Full) and NaturalReasoning) using the same Qwen3-235B-A22B-Thinking-2507-FP8 model.
As shown in Table 10, our Design-Logic-Reasoning datasets consistently outperform WebInstruct (Full) and NaturalReasoning across all benchmarks. Specifically, DLR-Book achieves the best results on MMLU, MMLU-Pro, and GPQA-Diamond, while DLR-Web attains the highest performance on SuperGPQA and the second-best on other benchmarks. These results validate the superior quality and effectiveness of our data generation approach compared to existing methods.
Table 10: Comparison with other doc-centric synthetic datasets on the Qwen3-8B-Base model. All experiments were conducted using a subsample of 304,181 instances from each dataset. The best results in each column are in bold, and the second-best are underlined.
Dataset MMLU MMLU-Pro GPQA-Diamond SuperGPQA
Pass@1 Pass@1 Pass@1 CoT-SC Pass@1
[…]
6.6 Ablation Studies
We conducted an ablation study on our question-synthesis pipeline to quantify the contributions of two key components: coarse-to-fine matching between design logics and corpus segments, and the use of explicit design logics. To ensure a fair comparison, all ablations synthesize questions from the same book corpus by uniformly sampling 304,181 text segments. Where applicable, we used identical retrieval with Qwen3-Embedding-4B and generated model responses with the same model, Qwen3-235B-A22B-Thinking-2507-FP8.
  • w/o Coarse Ranking: This setting bypasses the semantic similarity-based retrieval of the top-5 relevant design logics. Instead, the LLM is prompted to select the most suitable logic from a set of 5 randomly chosen logics.
  • w/o Fine Ranking: This setting removes the LLM-based re-ranking stage. The single most similar design logic (the top-1 result from the coarse retrieval) is used directly to generate the question.
  • w/ Example Questions: This setting replaces our use of abstract design logics. Instead of being guided by a logic, the LLM is prompted to generate a new question by imitating the style and structure of the most suitable exemplar. This exemplar is selected from a set of 5 relevant example questions that are retrieved from the question bank.
  • DESIGNER: This is our complete method, which includes coarse retrieval of the top-5 design logics by similarity, followed by LLM-based fine selection and synthesis based on “design logic”.
Table 11 summarizes the results of our ablation studies. Our complete “DESIGNER” approach consistently outperforms ablated configurations across most benchmarks. The removal of either coarse-grained retrieval (w/o Coarse Ranking) or fine-grained LLM re-ranking (w/o Fine Ranking) causes a noticeable performance drop, confirming the value of each stage in our two-step matching process. The “w/o Fine Ranking” configuration, in particular, performs well on the MATH benchmark, suggesting the initial retrieval is highly effective at identifying relevant logics. However, the consistent improvements across other benchmarks validate the full coarse-to-fine pipeline’s benefit for handling complex, multidisciplinary reasoning. Finally, the superior performance of design logics over concrete “Example Questions” confirms our hypothesis that explicit logical structures are more accurate and robust guides for high-quality question synthesis.
Table 11: Ablation study of our data synthesis pipeline. All experiments were conducted on the Qwen3-8B-Base model using data synthesized from the same book corpus. The best results in each column are in bold, and the second-best are underlined.
Method MATH-500 MMLU GPQA-Diamond SuperGPQA
Pass@1 CoT-SC Pass@1 Pass@1 CoT-SC Pass@1
[…]
6.7 Effect of Source Corpus Quality on Synthesized Data
Table 12: Effect of source-corpus quality on SFT outcomes with equal per-discipline and total size. All experiments were conducted on the Qwen3-8B-Base model. The best results in each column are in bold.
Source Corpus MATH-500 MMLU MMLU-Pro GPQA-Diamond SuperGPQA
Pass@1 CoT-SC Pass@1 Pass@1 Pass@1 CoT-SC Pass@1
[…]
It is widely accepted that book corpora are of higher quality than web corpora. To test how source corpus quality affects our data synthesis method, we conducted a controlled SFT experiment. We compared data synthesized from a high-quality book corpus to that from a web corpus. For a fair comparison, we ensured the disciplinary data distribution was identical across both datasets. For disciplines with limited data in the FineFineWeb corpus, we used all available instances. For other disciplines, we used random sampling. This process resulted in two final datasets of equal size (282,857 instances each) and identical per-discipline distribution.
As shown in Table 12, the results validate our hypothesis. The model fine-tuned on data from the book corpus consistently outperforms the one trained on data from the web corpus across almost all benchmarks. This performance gain is most notable in complex reasoning tasks like GPQA-Diamond and SuperGPQA, confirming that our synthesis method benefits from higher-quality source material. However, our method also proves robust to variations in source quality, as the performance gap between the two corpora is not substantial. This indicates that our approach effectively synthesizes high-quality questions to improve model performance, regardless of the initial corpus quality.
7 Related Work
Query-Centric Data Synthesis
The core idea of query-centric data synthesis is to generate new training data by iteratively expanding or evolving an existing query pool. Self-Instruct (Wang et al., 2023) samples questions from an existing query pool and leverages an LLM to generate new question-answer pairs, which are then used to fine-tune the model. Wizard LM (Xu et al., 2023) (Luo et al., 2025) (Luo et al., 2024) and Auto Evol-Instruct (Zeng et al., 2024) employ instruction evolution to produce more complex and diverse data. CoT-Self-Instruct (Yu et al., 2025) integrates the CoT reasoning mechanism into the instruction generation process, enhancing the quality of synthetic data through a “think-first, then-generate” approach. Prismatic Synthesis (Jung et al., 2025) pushes the definition of data diversity from superficial textual features to a more fundamental level of model behavior, aiming to maximize the diversity of model-induced gradients. The SPARQ method (Havrilla et al., 2025) introduces a framework for evaluating diversity and difficulty in data synthesis, arguing that high-difficulty data improves in-distribution performance while diversity enhances out-of-distribution generalization. Although these query-centric methods attempt to optimize the difficulty and diversity of synthetic data, they are still limited by the initial query pool and the distribution biases of the model used for synthesis, making it difficult to cover questions across various disciplines.
Doc-Centric Data Synthesis
Doc-centric data synthesis methods start with a large amount of unstructured (e.g., CC, books) or structured (e.g., knowledge graphs) textual data. They generate question-answer pairs by directly extracting or inferring them from documents, thereby ensuring the synthetic data is closely tied to specific domain knowledge and facts. UltraChat  (Ding et al., 2023) generates questions about world knowledge from knowledge sources like Wikidata and meta topics pre-defined by human experts (e.g., technology, health). Humpback (Li et al., 2024) also uses a large web corpus as a data source, training a “backward model” to infer an instruction that would lead to a given document. The KPDDS method (Huang et al., 2025) constructs a Topic-level Co-occurrence Probability Matrix (TCPM) based on extracted topics and knowledge points, guiding data synthesis by sampling from this matrix. MAmmoTH2 (Yue et al., 2024) mines high-quality questions directly from existing web content using a three-step “Recall-Extract-Refine” pipeline, but the difficulty of its extracted questions is low. NaturalReasoning (Yuan et al., 2025) also synthesizes data from a large-scale pre-training corpus, but its process is more focused on generating high-difficulty reasoning questions, though the diversity of its synthetic questions is limited by the single prompt. Our work uniquely introduces the concept of design logic, analogizing the doc-centric data synthesis process to a human teacher creating questions. By matching documents with appropriate design logic, we significantly increase the difficulty and diversity of the synthetic questions.
Reasoning Data Synthesis
Using synthetic long CoT data for data distillation is an effective way to improve the reasoning abilities of smaller models. DeepSeek (Guo et al., 2025) distills a series of smaller open-source models using high-quality data samples generated by DeepSeek-R1, significantly enhancing the reasoning capabilities of these smaller models. OpenMathReasoning (Moshkov et al., 2025) extracts high-quality math problems from an online math forum (AoPS) and leverages powerful existing models to generate long CoT responses for these problems, significantly improving the models’ mathematical abilities. OmniThought (Cai et al., 2025) collects questions from the fields of mathematics, code, and science, and generates a long CoT dataset based on multiple teacher models. However, these methods primarily focus on generating high-quality reasoning processes for existing questions, and the open-source community still lacks high-quality, multidisciplinary original questions. Our method, in contrast, focuses on leveraging naturally available multidisciplinary documents to synthesize difficult and diverse multidisciplinary questions by introducing design logic.
8 Conclusion
This paper introduces DESIGNER, a novel design-logic-guided data synthesis pipeline designed to address the scarcity of high-quality, multidisciplinary reasoning data for LLMs. Our core innovation is the concept of “Design Logic”, which abstracts the strategic process human experts use to create challenging questions. By leveraging these logics, we generated two large-scale datasets, DLR-Book and DLR-Web, comprising millions of complex questions across 75 disciplines from diverse raw text sources. Our synthesized questions are shown to be significantly more difficult and diverse than those from existing methods, moving beyond simple factual recall to require deep, multi-step reasoning. We validated the effectiveness of our approach through extensive experiments, demonstrating that models fine-tuned with our data achieve substantial performance gains in multidisciplinary reasoning. Specifically, models trained on our datasets not only outperform those fine-tuned with existing multidisciplinary datasets of the same volume but can also surpass the multidisciplinary reasoning performance of the official Qwen3 models of the same size, validating the effectiveness of our method. This work provides a scalable paradigm for creating diverse and challenging reasoning data, paving a new path for advancing LLMs’ reasoning capabilities beyond domain-specific tasks.
Appendix A Design Logic and Question Quantity
Table 13: Number of Design Logics, Book Questions, and Web Questions by Discipline
Discipline Design Logic Book Question Web Question
Aerospace Science and Technology 980 15025 1815
Agricultural Engineering 966 14950 11312
Agricultural Resources and Environment 409 14978 1443
Animal Husbandry 121 15023 318
Archaeology 776 15061 2572
Architecture 801 14931 4235
Art and Design 998 15164 12720
Astronomy 1600 40010 14915
Atmospheric Sciences 973 15029 3416
Basic Medicine 2289 49930 7118
Bioengineering 1041 11216 1059
Biology 4654 200078 111988
Biomedical Engineering 790 15120 5439
Business Administration 3873 99739 46337
Chemical Engineering and Technology 1393 20014 5235
Chemistry 6430 199839 68211
Chinese History 784 14876 14564
Chinese Language and Literature 998 14984 22233
Civil Engineering 950 14899 7773
Clinical Medicine 2844 99735 77182
Computer Science and Technology 4674 99796 219474
Control Science and Engineering 1338 20013 3254
Ecology 968 15113 4285
Economics 3864 99906 20064
Education 1287 20018 9824
Electrical Engineering 2904 59937 36949
Electronic Science and Technology 1445 19880 18864
English and Foreign Languages 985 29908 1344
Environmental Science and Engineering 1478 19954 28226
Ethnology 210 14856 51
Food Science and Engineering 768 15073 2470
Forensic Medicine 295 15010 34
Geography 2812 59987 12295
Geological Resources and Geological Engineering 173 15058 374
Geology 987 14958 5447
Geophysics 982 20050 3736
History of Science and Technology 494 14991 111
Hydraulic Engineering 664 14973 649
Information Resources Management 356 14968 69
Information and Communication Engineering 2367 39825 4777
Instrument Science and Technology 687 15017 378
Journalism and Communication 932 14847 4326
Law 2706 80110 62699
Management Science and Engineering 1317 20020 10440
Marine Sciences 852 14968 1228
Materials Science and Engineering 1949 40121 12263
Mathematics 9884 299464 181537
Mechanical Engineering 2955 59998 55356
Mechanics 1752 40046 4012
Mining Engineering 282 14986 438
Naval Architecture and Ocean Engineering 571 15034 515
Nuclear Science and Technology 1448 19988 1981
Nursing 767 15078 2228
Optical Engineering 927 14995 3272
Petroleum and Natural Gas Engineering 571 15040 988
Pharmacy 1829 39921 12756
Philosophy 4363 100029 128004
Physical Education 890 14912 10712
Physics 5768 199771 104982
Political Science 3268 59686 52586
Power Engineering and Engineering Thermophysics 704 14918 1199
Psychology 4336 99502 95585
Public Administration 960 14970 26669
Public Health and Preventive Medicine 1949 39820 19649
Remote Sensing Science and Technology 344 8081 271
Safety Science and Engineering 835 14986 409
Sociology 2093 39941 32361
Statistics 3388 59950 24806
Stomatology 592 15002 456
Surveying and Mapping Science and Technology 509 14964 358
Textile Science and Engineering 187 14978 325
Transportation Engineering 882 14929 4956
Urban and Rural Planning 576 15052 89
Veterinary Medicine 547 14918 3022
World History 987 39703 5503
Table 13: (Continued) Number of Design Logics, Book Questions, and Web Questions by Discipline
Appendix B Prompts
Prompt for Discipline Classification You are a professional multidisciplinary data labeling expert specializing in the classification of multidisciplinary academic questions. Please select the ONE most relevant label from the given list of discipline labels for the input question data. For question data that you cannot determine, use the ‘‘Unknown Discipline’’ label. Please directly output ‘‘labels’’: ‘‘(the label you selected)’’.
# List of Discipline Labels:
[‘Mathematics’, ‘Biology’, ‘Chemistry’, ‘Physics’, ‘Computer Science and Technology’, ‘Philosophy’, ‘Psychology’, ‘Business Administration’, ‘Clinical Medicine’, ‘Economics’, ‘Law’, ‘Political Science’, ‘Statistics’, ‘Electrical Engineering’, ‘Geography’, ‘Mechanical Engineering’, ‘Basic Medicine’, ‘Information and Communication Engineering’, ‘Sociology’, ‘Materials Science and Engineering’, ‘Pharmacy’, ‘Public Health and Preventive Medicine’, ‘Mechanics’, ‘Astronomy’, ‘World History’, ‘Bioengineering’, ‘English and Foreign Languages’, ‘Chemical Engineering and Technology’, ‘Electronic Science and Technology’, ‘Environmental Science and Engineering’, ‘Nuclear Science and Technology’, ‘Control Science and Engineering’, ‘Management Science and Engineering’, ‘Education’, ‘Geophysics’, ‘Art and Design’, ‘Agricultural Engineering’, ‘Aerospace Science and Technology’, ‘Atmospheric Sciences’, ‘Chinese Language and Literature’, ‘Civil Engineering’, ‘Ecology’, ‘Geology’, ‘Nursing’, ‘Optical Engineering’, ‘Public Administration’, ‘Journalism and Communication’, ‘Physical Education’, ‘Marine Sciences’, ‘Safety Science and Engineering’, ‘Architecture’, ‘Transportation Engineering’, ‘Power Engineering and Engineering Thermophysics’, ‘Food Science and Engineering’, ‘Archaeology’, ‘Biomedical Engineering’, ‘Chinese History’, ‘Veterinary Medicine’, ‘Instrument Science and Technology’, ’Hydraulic Engineering’, ‘Stomatology’, ‘Urban and Rural Planning’, ‘Petroleum and Natural Gas Engineering’, ‘Naval Architecture and Ocean Engineering’, ‘Surveying and Mapping Science and Technology’, ‘History of Science and Technology’, ‘Agricultural Resources and Environment’, ‘Remote Sensing Science and Technology’, ‘Information Resources Management’, ‘Mining Engineering’, ‘Forensic Medicine’, ‘Ethnology’, ‘Textile Science and Engineering’, ‘Geological Resources and Geological Engineering’, ‘Animal Husbandry’, ‘Other’, ‘Non-disciplinary’, ‘Unknown Discipline’]
# Example 1
Input: ‘‘Consider a photon traveling at the speed of light. How does the photon experience space, and what are the implications of relativistic beaming on its perception of spatial dimensions? Provide a detailed explanation, including any relevant mathematical derivations and physical principles.’’
Output: ‘‘labels’’: ‘‘Physics’’
# Example 2
Input: ‘‘A heavy pole, of mass M and length L, is freely hinged to a wall at the point O. A rope connects the other end of the pole, B, to a fixed point A on the wall above O. The system is in equilibrium, with the pole making an angle of $\theta$ with the horizontal, and the rope making an angle of $\alpha$ with the horizontal. Explore how the system’s parameters (M, L, $\theta$, $\alpha$) affect its equilibrium and stability.’’
Output: ‘‘labels’’: ‘‘Mechanics’’
# Example 3
Input: ‘‘If John rented a car for \$150 and had to buy 8 gallons of gas at \$3.50 per gallon to fill it up, and the final expense is \$0.50 per mile, how much did it cost him to drive 320 miles?’’
Output: ‘‘labels’’: ‘‘Mathematics’’
# Input Question Data
Input: ‘‘{text}’’
Output:
Figure 3: The few-shot prompt used for discipline classification. The model is presented with a list of disciplines and three examples, and is then asked to classify a given question, which replaces the {text} placeholder.
Prompt for Difficulty Classification You are an expert in education and examination, specializing in classifying the difficulty levels of multidisciplinary questions. For the given question, please evaluate its difficulty based on the complexity and length of the reasoning required to answer it. Label it as one of the following: **Easy**, **Medium**, **Hard**, or **Very Hard**. Please directly output ‘‘Difficulty: (Your chosen label)’’.
# Example 1
Input: ‘‘Consider a photon traveling at the speed of light. How does the photon experience space, and what are the implications of relativistic beaming on its perception of spatial dimensions? Provide a detailed explanation, including any relevant mathematical derivations and physical principles.’’
Output: ‘‘Difficulty: Very Hard’’
# Example 2
Input: ‘‘A heavy pole, of mass M and length L, is freely hinged to a wall at the point O. A rope connects the other end of the pole, B, to a fixed point A on the wall above O. The system is in equilibrium, with the pole making an angle of $\theta$ with the horizontal, and the rope making an angle of $\alpha$ with the horizontal. Explore how the system’s parameters (M, L, $\theta$, $\alpha$) affect its equilibrium and stability.’’
Output: ‘‘Difficulty: Hard’’
# Example 3
Input: ‘‘If John rented a car for \$150 and had to buy 8 gallons of gas at \$3.50 per gallon to fill it up, and the final expense is \$0.50 per mile, how much did it cost him to drive 320 miles?’’
Output: ‘‘Difficulty: Easy’’
# Given Question
Input: ‘‘{text}’’
Output:
Figure 4: The few-shot prompt used for difficulty classification. The model is presented with three examples and is then asked to classify the difficulty of a given question, which replaces the {text} placeholder.
Prompt for Question Type Classification You are an expert in education and examination, specializing in classifying question types. For the given question, please evaluate its question type and label it as one of the following: **Problem-solving question**, **Multiple-choice question**, **Proof question**, or **Other question types**. For any question that you cannot determine, use the ‘‘Other question types’’ label. Please directly output ‘‘Question type: (Your chosen label)’’.
# Example 1
Input: ‘‘Determine the number of $k$-letter sequences composed of the letters AA and $B$ such that the sequence contains at least two consecutive AA’s.’’
Output: ‘‘Question type: Problem-solving question’’
# Example 2
Input: ‘‘Consider the function $f(x)=\frac{e^{x}}{x}$. The value of the integral $I=\int_{1}^{\infty}\left(\frac{e^{x}}{x}-\frac{e^{-x}}{x}\right)dx$ is ___.’’
Output: ‘‘Question type: Other question types’’
# Example 3
Input: ‘‘Given that $a\in\{-1,2,\frac{1}{2},3,\frac{1}{3}\}$, if $f(x)=x^{a}$ is an odd function and is monotonically increasing on $(0,+\infty)$, then the possible values of the real number $a$ are ( ).
 A: 1,3-1,3
 B: $\frac{1}{3},3$
 C: $1,\frac{1}{3},3$
 D: $\frac{1}{3},\frac{1}{2},3$’’
Output: ‘‘Question type: Multiple-choice question’’
# Given Question
Input: ‘‘{text}’’
Output:
Figure 5: The few-shot prompt used for question type classification. The model is presented with three examples and is then asked to classify the type of a given question, which replaces the {text} placeholder.
Prompt for Reasoning-Oriented Filtering You will be provided with text from the internet.
Evaluate the following text extract for its potential usefulness for studying reasoning process. Use the following 5-point scoring system described below. Start from 0, points are accumulated based on the satisfaction of each criterion:
(1) Add 1 point if the extract contains any reasoning or thinking process.
(2) Add 1 point if the extract contains any explicit subgoal setting, where the writer breaks down the problem into smaller, intermediate goals. Subgoal setting might look like:
 • ‘‘First, we need to find ..., then we can determine ...’’
 • ‘‘To solve ..., let’s first ..., then ...’’
 • ‘‘Let’s tackle ... in three parts: (1) ..., (2) ..., and (3) ...’’
 • ‘‘To ..., I’$l$ first ..., then ...’’
(3) Add 1 point if the extract contains any verification steps. We want to mark instances where the writer explicitly checks their own work, such as by comparing the result to a known value or by checking the result of a calculation. Verification steps might look like:
 • ‘‘Let’s check ...’’
 • ‘‘To verify this is correct, I’$l$ ...’’
 • ‘‘Let’s test ... with a simple case: ...’’
 • ‘‘To ensure this solution is valid, I’$l$ check if ...’’
(4) Add 1 point if the text contains any backtracking behavior, where the writer realizes a path won’t work and explicitly goes back to try a different approach. An example of backtracking is: ‘‘Let me try again’’, ‘‘Wait’’, ‘‘I made a mistake’’, or ‘‘we need to try a different sequence of operations’’. We want to mark instances where the writer abandons a thought and backtracks to a previous computation.
(5) Add 1 point if the text contains any backward-chaining behavior, where the writer is working towards a goal but starts from the goal and works backward. It might like:
 • ‘‘To solve ..., let’s start with what we want to prove: ...Let’s verify this.’’
 • ‘‘If we want to find ..., let’s start with the desired result and work backward.’’
 • ‘‘To determine ..., I know the result ... Working backward from this final state using
# Task Format
Format your response in markdown as follows:
## Thoughts
[Brief description describing what behavior was noticed and where subgoal setting may have occurred, less than 100 words]
## Final score
[total points]
# Text to evaluate for reasoning degree
{text}
# Response
Figure 6: The prompt used for the reasoning-oriented filtering task. It defines a five-level scoring rubric to assess the usefulness of a text (which replaces the {text} placeholder) for studying reasoning.
Prompt for Design Logic Extraction You are an expert educator and a specialist in exam question design. Below, I have provided an exam question. Your task is to deduce the thought process of the question designer. Analyze how they constructed this question based on the relevant knowledge points. You need to go beyond the specific details of the question and its knowledge points to abstract and summarize the underlying design logic and principles behind the question.
The goal is for me to be able to use this abstracted design logic to create other high-quality, challenging questions that require complex logical reasoning for different knowledge points and source materials.
**Finally, you must organize the abstracted question-design logic you have summarized into English Mermaid format.**
--- Analyze the Question Design Logic from the Following Question ---
**Question:**
{text}
Figure 7: The prompt used for design logic extraction. The model is instructed to reverse-engineer the thought process behind a given question (which replaces the {text} placeholder) and to structure the abstracted logic in Mermaid format.
Instruction for Design Logic Retrieval Given a book snippet, retrieve the most suitable question-design logic in Mermaid format for creating a challenging exam question from the book snippet.
Figure 8: The task-specific instruction used for retrieving the most suitable design logic for a given text segment. Embeddings for both text segments and design logics are computed under this instruction using the Qwen3-Embedding-4B model, enabling similarity-based retrieval.
Prompt for Question Synthesis You are an expert in the field of education and examination design, and you are writing exam questions. Your task is to use the provided text to generate a high-quality exam question. Please follow the steps below to generate an English exam question and a reference answer:
**1. Create an Exam Question:**
- Based on the provided source text, write a challenging exam question at the graduate-level or above.
- Below are five question-design logics provided in Mermaid format. You need to select the most suitable question-design logic for creating a challenging question from the source text, and then strictly follow the corresponding question-design logic and steps to create a challenging question. Please record which design logic you used (by number) and output the corresponding numeric ID in the ‘‘id’’ field of the JSON below.
- The question should require critical thinking and test deep understanding and problem-solving skills, not just simple fact recall.
- The question must be self-contained and answerable without using the source text. If the question you write requires an answer based on the content of the source text, you must include the corresponding content and information from the source text within the question itself to make it self-contained.
- Ensure the question is self-contained, clear, without missing information or ambiguity, and has a correct answer.
- For multiple-choice questions, you should first analyze and determine the answer, then design the options to ensure that one specific option is the correct answer. The questions you design need to include as many options as possible (four or more). Do not be limited to only four options (A, B, C, D).
**2. Provide the Reference Answer:**
- Use the information in the source text to write a concise and accurate reference answer to the question you just created.
[…]
**At the end of your response, please organize your results into the following JSON format:**
‘‘exam_question’’: ‘‘*(Your question goes here)*’’,
‘‘reference_answer’’: ‘‘*(Your reference answer goes here)*’’,
‘‘id’’: ‘‘*(The ID of the logic you selected goes here)*’’
**--- Question-Design Logic 1 ---**
‘‘‘Mermaid
{logic1}
**--- Question-Design Logic 2 ---**
‘‘‘Mermaid
{logic2}
**--- Question-Design Logic 3 ---**
‘‘‘Mermaid
{logic3}
**--- Question-Design Logic 4 ---**
‘‘‘Mermaid
{logic4}
**--- Question-Design Logic 5 ---**
‘‘‘Mermaid
{logic5}
**--- Source Text for Question Creation ---**
{text}
Figure 9: The prompt used for question synthesis. The LLM is provided with a source text and five candidate design logics retrieved via semantic similarity. It is instructed to select the most suitable logic and strictly follow it to generate a graduate-level question and a corresponding reference answer, structured in a JSON format. The placeholders {logic1} through {logic5} and {text} are replaced with specific design logics and the source text, respectively.